TC = 100Q - 10Q^2 +5Q^3

To find optimal level of output
1
Expert's answer
2020-11-02T10:18:34-0500

The optimal level of output is output that maximizes a firm's profit. This level of output exists where marginal cost equals marginal revenue, that is, MC = MR.


Given the TC function, differential calculus is applied to find MC.


"MC = \dfrac {d}{dQ}(TC)"


"=\dfrac {d}{dQ}(100Q-10Q^2+5Q^3)"


"= 100-20Q+15Q^2"


Since, MR is not given, Let us Assume MR = $95.

"Note:" if MR is low, for example, "MR \leq \$90" , output cannot be found. Hence, MR = $95 is a better estimate.


The optimum output occurs where:

"MC = MR"

"100 - 20Q+15Q^2 =95" "=>15Q^2 -20Q+100 -95 =0"

"=>15Q^2-20Q+5 = 0"

Dividing by 5 each term gives:

"=> 3Q^2 - 4Q + 1 = 0"

"=> 3Q^2-3Q-Q+1=0"

"=> 3Q(Q-1)-1(Q-1)=0"

"=>(3Q-1)(Q-1)=0"

"\therefore \space Q = \dfrac {1}{3} \space or \space Q=1"


Thus, the optimal quantity, given that MR = $95, is 1 unit.